### v3

@@ 3.4.2 Hyperparameters
In all experiments, the stochastic gradient descent is chosen as the optimizer and the polynomial annealing policy [5] is used to schedule the learning rate. The hyperparameters and their values in this study are shown in Table 1.
Table 1: Hyperparameters
@@ 3.4.3 Performance metrics
Following [5, 6, 40, 7, 27], the mainstream performance metrics such as $Precision$, $Recall$, $F1score$, $Landslide\_IoU$ and mean Intersection over Union ($mIoU$) are employed, which are defined as follows
@@ 4 Results and Discussion
In this section, we carry out comparison experiments, ablation experiments, and cross validation experiments, along with discussions on the experimental results. To better illustrate the training process of HPCL-Net, it is shown in the form of pseudocode in Appendix.
@@ 4.1 Overall comparative experiments
Table 2: Numerical results of segmentation networks
@@
Figure 8: Graphical segmentation results on the relic landslides. From top to bottom are RGB, DEM, label, FFS-Net and ours
To visualize the landslide features learned by each model, the Gradient-weighted Class Activation Mapping (Grad-CAM) scheme [37] is employed to produce a heatmap, where features with more contribution to the target class have higher heat. We apply Grad-CAM to the last layer of the encoders of the FFS-Net and HPCL-Net, and the heatmaps are shown in Fig. 9. Compared to the FFS-Net, our model has more features with higher heat located in the back and side walls of the landslides, proving that the proposed model is capable of effectively learning the crucial features of relic landslides.
Figure 9: Visualization results for the last layer of the encoder by Grad-CAM. The landslides are outlined by the red line
@@ 4.2 Ablation experiments
Table 3: Numeric results of ablation experiments
@@
Figure 10: Visualization results for the encoder feature map: a label, b B, c B+F, d B+F+C, e B+F+C+G
@@
Figure 11: Visualization results for the last layer of the feature extraction and fusion module in FFS-Net and Heterogeneous Feature Extractor in HPCL-Net. The outline of the target landslide is highlighted in red line. a contour map, b B, c B+F
@@
Figure 12: Visualization results for the last layer of encoder by Grad-CAM. The outlines of landslides are highlighted by red line. a label, b B, c B+F, d B+F+C, e B+F+C+G
@@ 4.3 Cross validation experiments
Table 4: Cross validation of fusion method
@@
Table 5: Numeric results of two DEM distillation schemes
@@
Figure 13: Visualization results of two DEM distillation schemes
@@ 4.4 Complexity analysis
Table 6: Numerical results of Complexity
@@ 5 Conclusions
Appendix
@@
Figure 14: The landslide distribution maps from the six counties

### v4

@@ 3.4.2 Hyperparameters
In all experiments, the stochastic gradient descent is chosen as the optimizer and the polynomial annealing policy (Chen et al., 2017) is used to schedule the learning rate. The hyperparameters and their values in this study are shown in Table 1.
Table 1: Hyperparameters.
@@ 3.4.3 Performance metrics
Following Chen et al. (2017, 2018b); Wang et al. (2021); Chen et al. (2018c); Liu et al. (2023), the mainstream performance metrics such as $Precision$, $Recall$, $F1score$, $Landslide\_IoU$ and mean Intersection over Union ($mIoU$) are employed, which are defined as follows
@@ 4 Results and Discussion
In this section, we carry out comparison experiments, ablation experiments, and cross validation experiments, along with discussions on the experimental results. To better illustrate the training process of HPCL-Net, it is shown in the form of pseudocode in Appendix B.
@@ 4.1 Overall comparative experiments
Table 2: Numerical results of segmentation networks.
@@
Figure 8: Graphical segmentation results on the relic landslides. From top to bottom are RGB, DEM, label, FFS-Net and ours.
To visualize the landslide features learned by each model, the Gradient-weighted Class Activation Mapping (Grad-CAM) scheme (Selvaraju et al., 2017) is employed to produce a heatmap, where features with more contribution to the target class have higher heat. We apply Grad-CAM to the last layer of the encoders of the FFS-Net and HPCL-Net, and the heatmaps are shown in Fig. 9. Compared to the FFS-Net, our model has more features with higher heat located in the back and side walls of the landslides, proving that the proposed model is capable of effectively learning the crucial features of relic landslides.
Figure 9: Visualization results for the last layer of the encoder by Grad-CAM. The landslides are outlined by the red line.
@@ 4.2 Ablation experiments
Table 3: Numeric results of ablation experiments.
@@
Figure 10: Visualization results for the encoder feature map. (a) label, (b) B, (c) B+F, (d) B+F+C, (e) B+F+C+G.
@@
Figure 11: Visualization results for the last layer of the feature extraction and fusion module in FFS-Net and Heterogeneous Feature Extractor in HPCL-Net. The outline of the target landslide is highlighted in red line. (a) contour map, (b) B, (c) B+F.
@@
Figure 12: Visualization results for the last layer of encoder by Grad-CAM. The outlines of landslides are highlighted by red line. (a) label, (b) B, (c) B+F, (d) B+F+C, (e) B+F+C+G.
@@ 4.3 Cross validation experiments
Table 4: Cross validation of fusion method.
@@
Table 5: Numeric results of two DEM distillation schemes.
@@
Figure 13: Visualization results of two DEM distillation schemes.
@@ 4.4 Complexity analysis
Table 6: Numerical results of Complexity.
@@ 5 Conclusions
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The authors do not have permission to share data.
Acknowledgements and Funding
This work was supported by the National Key Research and Development Program of China [Grant No. 2021YFC3000400), and by the High Level Talent Team Project of the New Coast of Qingdao New District (Grant No. RCTD-JC-2019-06).
Appendix A Distribution Maps
@@
Figure 1: The landslide distribution maps from the six counties.
Appendix B Training process